Williamson Industries has $8 billion in sales and $1 billion in fixed assets. Currently, the company's fixed assets are operating at 95% of capacity. What level of sales could Williamson Industries have obtained if it had been operating at full capacity? Write out your answer completely. For example, 25 billion should be entered as 25,000,000,000. Round your answer to the nearest cent. $ What is Williamson's target fixed assets/sales ratio? Round your answer to two decimal places. % If Williamson's sales increase 12%, how large of an increase in fixed assets will the company need to meet its target fixed assets/sales ratio? Write out your answer completely. For example, 25 billion should be entered as 25,000,000,000. Round your answer to the nearest cent. Negative amount should be indicated by a minus sign. Do not round intermediate calculations. $
Answer #1

a). Full Capacity Sales = Current Sales(S0) / Current Operating Capacity

= $8,000,000,000 / 0.95 = $8,421,052,632

b). Target Fixed Assets/Sales Ratio = Fixed Assets / Full Capacity Sales

= $1,000,000,000 / $8,421,052,632 = 0.11875, or 11.88%

c). S1 = S0 * (1 + g) = $8,000,000,000 * (1 + 0.12) = $8,960,000,000

There will be no fixed assets will be required upto full capacity sales, i.e., $8,421,052,632

So, Increase in Fixed Assets = Target Fixed Assets/Sales Ratio * [S1 - Full Capacity Sales]

= 0.11875 * [$8,960,000,000 - $8,421,052,632]

= 0.11875 * $538,947,368.40 = $64,000,000